OpenStudy (dumbsearch2):

The directions on a box of muffin mix calls for 1-2/3 cup muffin mix and 4/3 cups water. If there is only 1 cup of muffin mix left in the box, how much water should be used?

OpenStudy (kropot72):

To reduce the 1-2/3 cups of muffin mix to 1 cup you divide 1-2/3 by 1-2/3. therefore the amount of water to go with 1 cup of muffin mix is found as follows: \[\frac{\frac{4}{3}}{\frac{5}{3}}=\frac{4}{3}\times \frac{3}{5}=you\ can\ calculate\]

OpenStudy (kropot72):

@dumbsearch2 Have you found the answer?

OpenStudy (dumbsearch2):

so it's 0.8?

OpenStudy (dumbsearch2):

thx i found it. :)

OpenStudy (dumbsearch2):

you solved it, thx

OpenStudy (kropot72):

You're welcome :)

OpenStudy (dumbsearch2):

your a GENIOUS no I'm serious. :) Like are u a teacher lol!

OpenStudy (kropot72):

Thank you. It's rewarding to have help appreciated.

OpenStudy (dumbsearch2):

yes! well do you get money for this?

OpenStudy (kropot72):

No, helping and learning is rewarding enough.

OpenStudy (dumbsearch2):

aw your amazing! :D

OpenStudy (kropot72):

Experience and practice is what it takes:)